OpenStudy (anonymous):

simplify the ratio 15 to 55

OpenStudy (anonymous):

divide by the GCD of 15 and 55

OpenStudy (anonymous):

This is the same as putting a fraction into lowest terms.\[15:55=\frac{15}{55}=\frac{3∗5}{11∗5}=\frac{3}{11}\]